# Chrono[Uncaptioned image]: A Simple Blueprint for Representing Time in MLLMs

### v5

@@ B.2 Implementation Details
In this section, we present the specific finetuning and evaluation details for Chrono-BLIP, as well as further specific zero-shot evaluation elements for Chrono-GPT
B.2.1 Training Details for BLIP2
@@
Considering the relatively small scale of the datasets, we choose not to finetune the entire 3 billion parameter-LLM backbone of our model but rather leverage parameter-efficient finetuning (Hu et al., 2022) to train only 19 million parameters, as described in Section 3.2.
@@
Since our objective is a generative language modeling task, our model can output minor formatting inconsistencies that would invalidate the prediction if not treated otherwise. Similar to (Yang et al., 2023), we post-process the model’s outputs by applying heuristics to clean up minor flaws in the prediction.
@@
For Charades-STA, we extract 20 video frames for each video, which is significantly less than prior work (Yan et al., 2023), which requires 128 frames. We train for up to 20 epochs with a batch size of 32 samples split across 4 A100-80GB GPUs for about 20 GPU hours. For QVHighlights and ActivityNet Captions, we extract 60 video frames for each video, which is less than required by prior work (Moon et al., 2023a, b). We train for up to 50 epochs with an effective batch size of 32 samples by splitting them across 8 A100-80-GB GPUs and leveraging accumulated gradients set to 4 (i.e., we have one sample per GPU) and train for about 170 GPU hours.
@@
We start with a learning rate (lr) of 1e-8, perform linear lr warmup to 3e-4 for 10% of the total number of iterations, and then apply a cosine lr decay. We utilize the AdamW (Loshchilov & Hutter, 2017) optimizer and sample the frames randomly during training.
B.2.2 Video processing
We build on top of the Salesforce lavis repository and leverage their implementation of randomly and uniformly extracting frames from a video. During training, we randomly sample frames as a means of data augmentation to alter the frames seen by the model in each batch. At inference, we sample uniformly to provide an equal coverage of the entire video. For random sampling, we start by uniformly extracting n+1 timestamps, where the first and last timestamps are at $t=0$ and $t=video\_length$, respectively. Afterward, we randomly sample one frame between each pair of adjacent timestamps, yielding the final n randomly sampled frames. This process of random sampling can be considered adding noise to a uniform sampling process.
B.2.3 Finetuning Hyperparameters
@@
Table 5: Hyperparameters for our base models for Charades-STA (Gao et al., 2017), QVHighlights (QVH) (Lei et al., 2021a), and ActivityNet (ANet) (Krishna et al., 2017) and for training of the answerer for NeXT-GQA (Xiao et al., 2023). LR: Learning rate.
@@
Table 5 shows a detailed list of hyperparameters. To ensure a fair comparison for all our models and ablations, we stick to hyperparameters presented in the table, except for the cases where we ablate certain parameters (Table 1). The only hyperparameters that change across benchmarks are the number of steps for the learning rate warmup and the number of input frames. The number of warmup steps depends on the dataset size and corresponds to 10% of the total number of steps, i.e.
#$\#steps_{warmup}=0.1*\#steps/epoch*\#epochs$ (3)
B.2.4 Zero-shot GPT-4o experiment details
@@
We query gpt-4o-2024-08-06 through the API. We use low image details for all queries. It is likely that with high image details, the results would be better, as more tokens are devoted per image, but we didn’t explore it in detail. We use temperature zero and maximum generation length for all the queries and use other API parameters as default. We observe some non-determinism even when using temperature 0, but all the results are statistically significant and, when possible, are averaged over at least two runs.
Appendix C Additional Ablations
C.1 Effect of number of trainable parameters
Table 6: Ablation on the LoRA Rank and number of trainable parameters.
Rank (# train. params) R1@0.5 R1@0.7 mIoU
2 (6M) 66.03 43.03 56.59
4 (10M) 65.71 43.59 56.78
8 (19M) 67.28 46.70 57.46
16 (37M) 67.72 45.38 57.35
In the main paper, we state the use of parameter-efficient finetuning techniques and thus training only a small number of parameters. In practice, we leverage LoRA (Hu et al., 2022) with a rank set to 8 and apply it to all linear layers in the large language model, yielding about 19 million trainable parameters.
In Table 6, we show the effect of different LoRA ranks and their resulting number of trainable parameters. We observe that setting the rank to 16, i.e. training double the number of parameters does not yield any significant improvement (row 4). Moreover, training fewer parameters, i.e. 10 or 6 million, by setting the rank to 4 or 2, respectively, does degrade the performance of Chrono (rows 1 and 2).
C.2 Timestamp design for GPT-4o
Table 7: Ablation studies on Charades-STA Validation set. Effect of timestamp design on zero-shot evaluation using GPT-4o.
# Rep. Prec. Inter. R1@0.5 R1@0.7 mIoU
(1) Rel Dec 32.85 11.38 37.78
(2) Abs Dec 32.38 13.30 36.99
(3) Rel Int 34.68 15.46 38.04
(4) Abs Int 29.91 11.28 33.43
(5) Rel Int 34.86 11.20 38.44
(6) Abs Int 28.35 11.29 35.43
(7) Abs Dec 28.72 10.74 34.78
Table 8: Ablation studies on QVHighlights Validation set. Effect of timestamp design on zero-shot evaluation using GPT-4o.
# Rep. Prec. Inter. R1@0.5 R1@0.7 mIoU
(1) Rel Dec 48.39 26.06 46.45
(2) Abs Dec 44.19 24.58 43.82
(3) Rel Int 30.42 14.90 33.91
(4) Abs Int 36.77 20.68 37.41
(5) Rel Int 62.84 42.19 57.91
(6) Abs Int 61.68 41.80 57.12
(7) Abs Dec 58.77 40.16 55.83
Tables 7 and 8 show our ablations on zero-shot prompting GPT-4o (using the prompt shown in Figure 4) with different timestamp representations, precision, and whether we interleave them with the frames or not. Table 8 supports our main hypothesis, that interleaving timestamps between frames leads to better localisation, also in the zero-shot case.
The effect of interleaving is less clear with Charades-STA, as shown in Table 7. We hypothesize that this difference is caused by the fact that we use only 20 frames (out of shorter videos) for Charades-STA, whereas in the QVHighlights, GPT-4o observes 60 frames (of videos up to several minutes long). This means appended timestamps are further away from the frame that they refer to, making the gap larger to interleaving frame timestamps in Table 8. However, in Charades-STA, the shorter distance between token frames and appended timestamps seemingly has a smaller effect. Additionally, we could hypothesize that successive frames are more in distribution for GPT-4o than a series of frames with interleaved timestamps, therefore offering some benefit to simply appending the timestamps in very short frame sequences. This illustrates how some timestamp decisions that are easier to learn and used if finetuned on, as shown for Chrono, might be more out-of-distribution and hence less performant when used with models zero-shot. To a smaller degree, we can also observe a similar trend regarding relative versus absolute timestamps. The preference for the former is higher in shorter videos but still maintains for longer ones. We hypothesize that this better performance of relative timestamps in Charades-STA might be caused by relative timestamps offering higher temporal resolution in shorter videos. Ultimately, although absolute timestamps might be better when finetuned, as we showed for Chrono, we find that relative timestamps are comparable or better when using GPT-4o zero-shot. Lastly, the best configurations for both datasets use integer precision rather than decimal, even if the latter offers a higher temporal resolution, which is in agreement with the results when finetuning, as shown in Table 1.
C.3 Chrono design for NExT-GQA
The best Chrono-GPT setup for NExT-GQA uses a single stage, i.e., asks for grounding and question answering in the same model response. We find that this performs better than 2 stages, where the model first performs moment retrieval, and then the model is asked to answer the question. This is the case even when the second stage is able to zoom in on the window, use more frames overall, and keep the first stage in context. We choose to use 1.4FPS for the final results. This uses an equivalent number of frames across the entire dataset (with an average length of around 40 seconds) compared to using 60 frames for every video. Using absolute interleaved timestamps significantly improves temporal grounding and, therefore, grounded accuracy. It modestly helps overall accuracy, i.e. , not thresholded by grounding, as can be seen also in Table 9. This could indicate that including temporal grounding capabilities into models that process video cannot only offer greater interpretability but also improve their overall performance at other tasks like video QA.
Table 9: Does moment retrieval help Video QA? We randomly choose 1000 samples from NExT-GQA validation split. We test whether asking the model to perform moment retrieval simultaneously to answering the question improves accuracy. We test whether timestamping the frames has an effect on this. Chrono-GPT uses interleaved timestamps with absolute time. All models use a single stage with 1.4FPS.
MR Time mIoU mIoP IoP@.5 A@GQA A@QA
- - - - 79.80
22.47 29.85 27.13 22.03 80.07
37.02 51.62 52.45 43.35 80.15
@@ C.3 Chrono design for NExT-GQA
Given the video and the query, find the relevant windows. Think step by step. Reason about the events in the video and how they relate to the query. After your reasoning, output ‘ANSWER: $<$your answer$>$‘ in the format specified in the task prompt. Always provide a non-empty answer after your thoughts. If you think the event does not take place in the video, give your best guess, as otherwise the evaluation will be marked as incorrect. Never provide an empty list for $<$your answer$>$. The descriptions of moments are sometimes imprecise, so retrieve the closest moment. If you don’t see an event remotely similar to the description, guess what is the most likely moment given the context. For instance, for cutting onion this could be between the time we see that the scene takes place in the kitchen and the time we see the onions being boiled in the pan. The answer should be in the format of a list indicating the start and end of a window of moment, [start_window, end_window], for instance [0, 1]. If you detect multiple windows for the same moment, choose the most relevant one. It’s important your final answer only contains one window. It is very important that the answer is in this format, otherwise the evaluation will fail.
@@
Appendix D Additional Qualitative Results
In this section, we provide further qualitative results and discuss shortcomings of our model that can be addressed in future research. For each example, we provide the discussion in the respective caption for the convenience of the reader. Examples 1 through 6 are from QVHighlights (Lei et al., 2021a) and examples 6 through 10 are from Charades-STA (Gao et al., 2017).
@@ Appendix D Additional Qualitative Results
Figure 5: We observe the capability of our model to recognize and differentiate 3 separate moments that repeatedly occur with intermittent interruptions. Each predicted moment depicts the respective natural language query.
@@
Figure 7: Our model predicts a third window, that does not align with a ground truth moment, which, in fact, is accurate and depicts the query.
Figure 8: The video contains two groups of relevant moments which contain a short sub-2-second cut that does not depict the query. With a resolution of 60 frames for a 150-second long video, corresponding to a frame being seen every 2.5 seconds, Chrono-BLIP can not detect the cut and predicts two long moments that encompass the two short ones. Moreover, our model predicts a third window (red) that does not match a ground truth label. The predicted moment does in actuality depict soldiers, but those are not escorting other people.
Figure 9: Our model has trouble predicting very short moments in long videos and high-frequency jump cuts given the resolution of the frame sampling.
Figure 10: Given a very ambiguous query, our model’s prediction aligns with the ground truth.
Figure 11: Our model fails to recognize the action of the man closing the front door. We hypothesize this is because the door is very dark and poorly visible. Our model predicts the moment when the door is clearly visible and closed.
Figure 12: Our model fails to recognize the action of putting down the broom at the beginning of the video and predicts the moment when the man picks the broom back up and then holds it.

### v6

@@ Appendix A Implementation Details
In this section, we present the specific finetuning and evaluation details for Chrono-BLIP and Chrono-Qwen, as well as further specific zero-shot evaluation elements for Chrono-GPT
A.1 Finetuning Details for BLIP-2 and Qwen2.5-VL
@@
Considering the relatively small scale of the datasets, we choose not to finetune the entire model but rather leverage parameter-efficient finetuning [hu2021lora]. For the BLIP-2 LLM backbone, this amounts to training only 19 million parameters, instead of 3 billion as described in Section 3.2.
@@
Since the objective is a generative language modeling task, the model can output minor formatting inconsistencies that would invalidate the prediction if not treated otherwise. Similar to [yang2023vid2seq], we post-process the model’s outputs by applying heuristics to clean up minor flaws in the prediction.
@@
For Charades-STA, we extract 20 video frames for each video, which is significantly less than prior work [yan2023unloc], which requires 128 frames. We train for up to 20 epochs with a batch size of 32 samples, using A100-80GB GPUs for training. For BLIP-2, this takes about 20 hours when using data parallelism split across 4 A100-80GB GPUs. For QVHighlights and ActivityNet Captions, we extract 60 video frames for each video, which is less than required by prior work [moon2023CQ_DETR, moon2023QD-DETR]. We train for up to 50 epochs with an effective batch size of 32 samples. For BLIP-2, finetuning in a single 8 GPU node with gradient accumulation takes about 170 GPU hours.
@@
We start with a learning rate (lr) of 1e-8, perform linear lr warmup to 3e-4 for 10% of the total number of iterations, and then apply a cosine lr decay. We utilize the AdamW [loshchilov2019adamw] optimizer and sample the frames randomly during training.
Table IX shows a detailed list of hyperparameters. To ensure a fair comparison for all models and ablations, we stick to hyperparameters presented in the table, except for the cases where we ablate certain parameters (Table IV). The only hyperparameters that change across benchmarks are the number of steps for the learning rate warmup and the number of input frames. The number of warmup steps depends on the dataset size and corresponds to 10% of the total number of steps, i.e.
$steps_{warmup}=0.1*\#steps/epoch*\#epochs$ (3)
@@
TABLE IX: Hyperparameters for the base models for Charades-STA [gao2017charadesSTA], QVHighlights (QVH) [lei2021qvhighlights], and ActivityNet (ANet) [krishna2017ActivityNetCaptions] and for training of the answerer for NExT-GQA [Xiao2023NExT-QGA]. LR: Learning rate.
@@
A.2 Zero-shot experiment details
We query gpt-4o-2024-08-06 through the API. We use low image details for all queries.
@@ A.2 Zero-shot experiment details
We use temperature zero and maximum generation length for all the queries and use other API parameters as default. We observe some non-determinism even when using temperature 0, therefore all GPT-4o results are averaged over at least two runs.
A.3 Video processing
We build on top of the Salesforce lavis repository and leverage their implementation of randomly and uniformly extracting frames from a video. During training, we randomly sample frames as a means of data augmentation to alter the frames seen by the model in each batch. At inference, we sample uniformly to provide an equal coverage of the entire video. For random sampling, we start by uniformly extracting n+1 timestamps, where the first and last timestamps are at $t=0$ and t=video_lengtht=video\_length, respectively. Afterward, we randomly sample one frame between each pair of adjacent timestamps, yielding the final n randomly sampled frames. This process of random sampling can be considered adding noise to a uniform sampling process.
Appendix B Zero-shot Chrono Prompt for Zero-shot moment retrieval with GPT-4o
In Figure 4, we show the prompt used for querying GPT-4o in the zero-shot setting. This includes indications on the expected format, as well as encouragement to reason step by step before providing the final window.
@@ Appendix B Zero-shot Chrono Prompt for Zero-shot moment retrieval with GPT-4o
Given the video and the query, find the relevant windows. Think step by step. Reason about the events in the video and how they relate to the query. After your reasoning, output ‘ANSWER: <<your answer>>‘ in the format specified in the task prompt. Always provide a non-empty answer after your thoughts. If you think the event does not take place in the video, give your best guess, as otherwise the evaluation will be marked as incorrect. Never provide an empty list for <<your answer>>. The descriptions of moments are sometimes imprecise, so retrieve the closest moment. If you don’t see an event remotely similar to the description, guess what is the most likely moment given the context. For instance, for cutting onion this could be between the time we see that the scene takes place in the kitchen and the time we see the onions being boiled in the pan. The answer should be in the format of a list indicating the start and end of a window of moment, [start_window, end_window], for instance [0, 1]. If you detect multiple windows for the same moment, choose the most relevant one. It’s important your final answer only contains one window. It is very important that the answer is in this format, otherwise the evaluation will fail.
@@
Appendix C Additional Qualitative Results
In this section, we provide further qualitative results and discuss shortcomings of the proposed approach which can be addressed in future research. For each example, we provide the discussion in the respective caption for the convenience of the reader. Examples 1 through 6 are from QVHighlights [lei2021qvhighlights] and examples 6 through 10 are from Charades-STA [gao2017charadesSTA].
@@ Appendix C Additional Qualitative Results
Figure 5: We observe the capability of the model using the ChronoRefer to caption blueprint to recognize and differentiate 3 separate moments that repeatedly occur with intermittent interruptions. Each predicted moment depicts the respective natural language query.
@@
Figure 7: The model using the ChronoRefer to caption blueprint predicts a third window, that does not align with a ground truth moment, which, in fact, is accurate and depicts the query.
Figure 8: The video contains two groups of relevant moments which contain a short sub-2-second cut that does not depict the query. With a resolution of 60 frames for a 150-second long video, corresponding to a frame being seen every 2.5 seconds, Chrono-BLIP can not detect the cut and predicts two long moments that encompass the two short ones. Moreover, the model using the ChronoRefer to caption blueprint predicts a third window (red) that does not match a ground truth label. The predicted moment does in actuality depict soldiers, but those are not escorting other people.
Figure 9: The model using the ChronoRefer to caption blueprint has trouble predicting very short moments in long videos and high-frequency jump cuts given the resolution of the frame sampling.
Figure 10: Given a very ambiguous query, the prediction of the model using the ChronoRefer to caption blueprint aligns with the ground truth.
Figure 11: The model using the ChronoRefer to caption blueprint fails to recognize the action of the man closing the front door. We hypothesize this is because the door is very dark and poorly visible. The model using the ChronoRefer to caption blueprint predicts the moment when the door is clearly visible and closed.
Figure 12: The model using the ChronoRefer to caption blueprint fails to recognize the action of putting down the broom at the beginning of the video and predicts the moment when the man picks the broom back up and then holds it.